Problem 1 Do this problem by hand - you may use Excel for calculations The tensile strength of Portland cement is being studied. Several different mixing techniques can be used economically. A completely randomized experiment is conducted to determine if mixing technique has a significant effect on the mean tensile strength of cement. Several measurements of tensile strength are recorded (see table below) for each mixing technique Tensile Strength (lb/in2) measurements 1 324 448 364 654 4 504 485 422 693 Mixing Technique 461 467 465 537 402 600 502 655 398 369 417 538 4 (a) What is the response? What is the factor? How many levels of the factor are used? How many replicates were used? 0.05. Do (b) Test the hypothesis that mixing techniques affect the mean tensile strength of the cement. Use α mixing techniques affect the mean tensile strength of the cement? Yes or No?
Answer #1

1.Here the response is tensile strength measurements.

factor is mixing technique

there are four levels of factor.

there are no replications.

null hypothesis,H0:tow1=tow2=tow3=tow4 i.e,the treatment or factor effects are same.

alternative hypothesis,H1:At least two of towi's are different,i=1(1)4.

Raw S.S(R.S.S)=(324^2)+(461^2)+...+(538^2).=4909201

Grand Total(G)=324+...+538=9705

Correction Factor(C.F)=(G^2)/n=(G^2)/20=4709351.25,where n is total number of plots=4*5=20

Total S.S(T.S.S)=R.S.S-C.F=199849.75

Treatment S.S.=(((T1^2)+...+(T4^2))/5)-C.F=(24152911/5)-4709351.25=121230.95

[T1,T2,T3,T4 are treatment totals of treatment 1,2,3,4,such that

T1=2089,T2=2369,T3=2170,T4=3079]

Error S.S=Total S.S-Treatment S.S=78618.8

ANOVA TABLE

SOURCE

OF VARIATION

S.S d.f M.S.S=S.S/d.f variation ratio,'F'
TREATMENTS 121230.95 3 40410.32
ERROR 78618.8 16 4913.68
TOTAL 199849.75 19 45324

VARIATION RATIO 'F'=40410.32/4913.68=8.22

from f- distribution table,F0.05(3,16)=3.06

THUS 'F'>F0.05(3,16),THUS H0 is rejectedat 5%level of significance

thus mixing techniques affect the mean tensile strength of cement